OpenStudy (17jeisaac):

In a random sample of 70 people, it was found that 44 of them were fans of the New York Yankees. What is the margin of error for the true proportion of all individuals who are fans of the New York Yankees?

OpenStudy (perl):

We can use a formula. Margin of error = significance level coefficient*Standard error of p We are not given the confidence level, But if we assume a coefficient of 3 for 99.73% confidence level Standard error of p = sqrt [p*(1-p)/n] p = x/n = 44/70 = 0.6286 Therefore, Margin of error = 3 * sqrt [0.6286*0.3714/70]